# AppApp Readings

### v1

@@ Abstract
The frequency dependencies of the in-plane temporal spin stiffness (A10) for two electron cocentrations are presented in Fig. S4. Similarly to the out-of-plane susceptibility, discussed in the main text, it decays with the increase of frequency. At low temperatures the frequency spacing becomes finer, enabling accurate extrapolation to obtain the static limit. We note that while despite strong frequency dependence, $\chi_{\rm{op},\omega_{n}}$ is continuous at zero frequency, i.e. its zero-frequency value coincides with the considered static limit, the in-plane mode, being related to zero momentum transfer in the local coordinate frame, it however shows a jump at zeroth Matsubara frequency. The magnitude of the jump is higher at higher doping levels. Paradoxically, absence of the jump in the out-of-plane component yields its faster decay with frequency, since in that case the jump is replaced by smooth, yet rapidly decreased, function.
@@
Figure S4: Dynamical temporal spin stiffness for in-plane mode as a function of bosonic frequency $\omega_{n}$ for various temperatures and electron concentrations.
.6 The details of the comparison of the spin gap in different approaches at various temperatures
@@
Figure S5: The comparison of different approximations for spin gap determined from the Eq (4) of the main text at various temperatures. The dependence of energy gap logarithm $\log_{10}\Delta$ on electron concentration $n$ is computed with inclusion of only zeroth bosonic frequency $\omega_{n}=0$, all frequencies but with temporal spin stiffness approximated by its static limit $\lim_{\omega_{n}\to 0}\chi_{\rm{op},\omega_{n}}$ and the calculation with dynamic temporal spin stiffness $\chi_{\rm{op},\omega_{n}}$ obtained as a function of bosonic frequency $\omega_{n}$.
@@
In Fig. S5 we present the dependence of the magnetic energy gap $\Delta$ (in logarithmic scale) on electron concentration $n$ for various temperatures. The results are obtained from equation (3) of the paper using three different approaches: (i) including only the zero-frequency contribution $\omega_{n}=0$, (ii) approximating the dynamic temporal spin stiffness $\chi_{\text{op},\omega_{n}}$ by its static limit $\lim\limits_{\omega_{n}\to 0}\chi_{\text{op},\omega_{n}}$, and (iii) using the exact values of $\chi_{\text{op},\omega_{n}}$ at all bosonic Matsubara frequencies $\omega_{n}$. It can be seen that both approaches (i) and (ii) results in exponential decrease of the magnetic energy gap with lowering temperature, although inclusion of finite bosonic frequencies increases $\Delta$ by several orders of magnitude. Full inclusion of temporal spin stiffness frequency dependence results in a completely different trend, that is described in the main text: only for states in a certain region of electron concentrations $n$ around $n\approx 0.9$ does the magnetic energy gap significantly decrease with lowering temperature while approaching a certain limit outside of this stability region.

### v3

@@ 6 Frequency dependence of the in-plane temporal spin stiffness
The frequency dependencies of the in-plane temporal spin stiffness (S20) for two electron concentrations are presented in Fig. 4. Similarly to the out-of-plane susceptibility, discussed in the main text, it decays with the increase of frequency. At low temperatures the frequency spacing becomes finer, enabling accurate extrapolation to obtain the static limit. We note that while despite strong frequency dependence, $\chi_{\rm{op},\omega_{n}}$ is continuous at zero frequency for states with spiral magnetic order, i.e. its zero-frequency value coincides with the considered static limit. However, the in-plane mode, corresponding to zero momentum transfer in the local coordinate frame, exhibits a discontinuity at zero Matsubara frequency. This discontinuity reflects the difference between the dynamical limit of the retarded susceptibility, $\chi(\omega)$, and the static thermal spin susceptibility, $\chi^{\text{th}}=\chi({\rm i}\omega_{n}=0)$. Physically, this difference appears due to conservation of total magnetic moment and the ergodicity of the system’s state, similarly to previous study of the paramagnetic state near Mott transition FrequencyAndJumpIssues . The magnitude of the jump is higher at higher doping levels. Paradoxically, absence of the jump in the out-of-plane component yields its faster decay with frequency, since in that case the jump is replaced by smooth, yet rapidly decreased, function. From the ergodicity point of view, the infinite-time correlations related to the jump of the in-plane temporal spin stiffness are replaced by a large but finite correlation time of the out-of-plane correlations, manifested as a sharp peak at $\omega=0$ with finite width.
@@
Figure 4: Dynamical temporal spin stiffness for in-plane mode as a function of bosonic frequency $\omega_{n}$ for various temperatures and electron concentrations.
@@
Figure 5: The comparison of different approximations for spin gap determined from the Eq (3) of the main text at various temperatures. The dependence of energy gap logarithm $\log_{10}\Delta$ on electron concentration $n$ is computed with inclusion of only zeroth bosonic frequency $\omega_{n}=0$, all frequencies but with temporal spin stiffness approximated by its static limit $\lim_{\omega_{n}\to 0}\chi_{\rm{op},\omega_{n}}$ and the calculation with dynamic temporal spin stiffness $\chi_{\rm{op},\omega_{n}}$ obtained as a function of bosonic frequency $\omega_{n}$.
7 The details of the comparison of the spin gap in different approaches at various temperatures
@@
In Fig. 5 we present the dependence of the magnetic energy gap $\Delta$ (in logarithmic scale) on electron concentration $n$ for various temperatures. The results are obtained from equation (2) of the main text using three different approaches: (i) including only the zero-frequency contribution $\omega_{n}=0$, (ii) approximating the dynamic temporal spin stiffness $\chi_{\text{op},\omega_{n}}$ by its static limit $\lim\limits_{\omega_{n}\to 0}\chi_{\text{op},\omega_{n}}$, and (iii) using the exact values of $\chi_{\text{op},\omega_{n}}$ at all bosonic Matsubara frequencies $\omega_{n}$. It can be seen that both approaches (i) and (ii) results in exponential decrease of the magnetic energy gap with lowering temperature, although inclusion of finite bosonic frequencies increases $\Delta$ by several orders of magnitude. Full inclusion of temporal spin stiffness frequency dependence results in a completely different trend, that is described in the main text: only for states in a certain region of electron concentrations $n$ around $n\approx 0.9$ does the magnetic energy gap significantly decrease with lowering temperature while approaching a certain limit outside of this stability region.